# Recent Advances in Attack and Defense Approaches of Large Language Models

### v2

@@ 3.4 Energy-Latency Attacks and Potential Threats against LLMs
The energy-latency attacks are originated in a broader area of neural networks and their inference processes. The NMTSloth methodology, as discussed in Feng et al. (2024), introduces a gradient-guided approach to detect efficiency degradation in Neural Machine Translation (NMT) systems. By delaying the appearance of the end-of-sequence (EOS) token through subtle perturbations, NMTSloth demonstrates how altering the output probability distribution can escalate computational demands. Another empirical study (Hong et al., 2021) leverages the adaptive nature of neural networks by introducing subtle perturbations during inference, significantly increasing the model’s inference time.
@@
LLMs, akin to other neural networks, are vulnerable to energy-latency attacks. Methods explored in the aforementioned paper can be similarly adapted for LLMs to heighten computational requirements and prolong response times. According to our research, we have identified that LLMs (LLaMa-series) tend to engage in excessive analysis in some trigger scenario’. For example, when the input data contains numerical values, using the instruction "let’s think step-by-step" prompts the model to prioritize mathematical computations, even if the instruction itself is not directly related to solving math problems. This tendency can lead the model to engage in unnecessary and detailed mathematical analysis, which may not align with the user’s intended query or task.
@@
Although energy-latency attacks have not been extensively explored in the context of LLMs, addressing their potential threats is crucial for ensuring the efficient and reliable operation of these models. While building efficient LLMs has been the focus of extensive research (Yang et al., 2024b; Sanh et al., 2020; Jiao et al., 2020; Cho, 2014), understanding and mitigating the risks posed by energy-latency attacks remains an important area for future investigation. Enhancing our ability to counteract these attacks will contribute to the development of more resilient and effective LLM systems.
@@ 4 Defense
This section explores advanced defense strategies aimed at enhancing the robustness and safety of LLMs. We categorize them into three main subsections: robustness enhancements, post-safety alignment, and model merge (Yang et al., 2024a) techniques. Robustness enhancements encompass proactive measures to strengthen LLMs against adversarial inputs, data biases, and other potential weaknesses. These include Red Teaming, Adversarial Training, Safety Fine Tuning, Post-Safety Alignment, and Model Merging. Post-safety alignment strategies focus on ensuring the outputs of LLMs align with the ethical standards and societal expectations, addressing issues such as fairness, transparency, and accountability. Model merge techniques involve integrating multiple models or methodologies to harness their combined strengths, enhancing overall performance, reliability, and adaptability of LLMs. By comprehensively reviewing these defense strategies, this section aims to provide insights into cutting-edge approaches for defending LLMs, ensuring their secure and effective deployment in real-world applications. Understanding and advancing these defenses are crucial steps toward building trustworthiness and resilience in LLMs, thereby facilitating their responsible integration into various domains of modern society.
@@ 4.1 Red Team Defense
Another effective way to discover potential risks in LLMs deployment is to use the Red Team Defense. The Red Team Defense methodology is centered around simulating real-world attack scenarios to uncover potential vulnerabilities and weaknesses in LLMs. The outcomes are used to improve security policies, procedures, and technical defenses. The process includes the following steps:
@@
Pain Points: Red teaming can be resource-intensive and requires skilled personnel to effectively mimic sophisticated attack strategies (Xu et al., 2021; Ribeiro et al., 2020; Röttger et al., 2021). The development of red teaming methodologies is still in its early stages, resulting in a lack of comprehensive statistical data on its effectiveness and outcomes.
Recent advancements in leveraging language models for red teaming have introduced automated approaches to simulate test cases and then employ a classifier to detect whether these test cases could lead to harmful outputs from the LLM Perez et al. (2022). The shift towards automation brings several benefits: firstly, it increases test case diversity and difficulty; secondly, it enables scalable testing across a wide range of scenarios and environments. The work proposed by Ganguli et al. (2022) aimed to build a more efficient AI-assistant interface to collect scale red team data for further analysis. They also studied the scalability of different sizes and types of LLMs under read team attacks, pointing out that rejection sampling and RLHF could build a stronger defense against various attacks.
Besides security and defense purpose, red teaming could help boost LLM performance in different areas as well. In the paper Buszydlik et al. (2023), the authors use red teaming techniques to simulate different types of mathematical problems and puzzles, and then evaluate the performance of LLMs in solving them. While the paper’s focus on mathematical tasks is narrow, its methodology and findings are valuable for understanding the broader impact of red teaming techniques on LLM performance.
@@ 4.2 Adversarial Training
Pain Point: Given the vast input space, comprehensively identifying failure modes in LLMs is challenging and resource-intensive. Besides, defenders often focus on robustness-related failure modes, such as those involving Lp-norm attacks. However, attackers may employ more subtle methods, including Trojans and jailbreaks, which are harder to detect. And adversarial training often causes a trade-off between robustness and performance on clean data (Min et al., 2020; Raghunathan et al., 2020). Furthermore, LLMs’ responses to adversarial training are less effective than expected. They might memorize adversarial samples used in training rather than developing generalized defenses. And they struggle to modify pre-train gained knowledge (Prakash et al., 2024b)
@@
Recent studies have introduced new approaches to address these challenges. Liu et al. (2020) find that adversarial pre-training using ALUM (Adversarial Learning with Unlabeled Model) leads to improvements in both generalization and robustness across a range of natural language processing tasks. The use of virtual adversarial training objectives (Miyato et al., 2018) allows ALUM to smooth embedding space of model and balance standard error with robust error effectively. This capability suggests a promising direction for adversarial training in LLMs, integrating robust defense strategies without compromising model performance. To further address the computational inefficiencies of traditional adversarial training methods, Xhonneux et al. (2024) introduces two novel algorithms that perform adversarial attacks within the continuous embedding space of LLMs. These algorithms—CAT (Continuous Adversarial Training) and CAPO (Continuous Adversarial Perturbation Optimization)—significantly enhance model robustness against discrete adversarial attacks while maintaining utility. By operating in the continuous embedding space, these algorithms reduce computational overhead and provide scalable solutions for adversarial training. In order to improve the generalization of adversarial training, Casper et al. (2024) employ latent adversarial training instead of generating adversarial examples aimed at specific failure modes. Their primary objective is to enhance the model’s robustness against potential shifts in data distribution that may occur between the model’s development and deployment phases in real-world applications. These shifts could include the introduction of Trojans, jailbreaks, and other unforeseen challenges.
@@
Future Work: Building on the insights from Casper et al. (2024), future research could explore the implementation of latent adversarial training as an alternative to traditional adversarial training during the fine-tuning stage. Additionally, incorporating adversarial training into the pre-training phase of model development could further enhance robustness against attacks. These approaches aim to improve the model’s resilience by addressing vulnerabilities early in the training process and better equipping the model to handle adversarial inputs.
@@ 4.3 Safety Fine-Tuning
Fine-tuning has gained great popularity among end-users as a core technique for customizing pre-trained LLMs to various downstream tasks. However, recent research has revealed potential safety risks associated with exercising this privilege. Qi et al. (2023b) find that the safety alignment of LLMs can be compromised by fine-tuning with only a few adversarially designed training examples. In addition, they further suggest that even without malicious intent, simply fine-tuning with benign and commonly used datasets may also inadvertently degrade the safety alignment of LLMs.
To address the vulnerabilities introduced by fine-tuning, a natural solution one would expect is to incorporate safety-related examples during the fine-tuning stage. Bianchi et al. (2023) and Zhao et al. (2024a) prove the effectiveness of this approach. And they suggest that including a small number of safety examples in the fine-tuning process significantly enhances the safety of LLMs while not degrading their usefulness. However, adding excessive safety samples may lead to LLMs rejecting safe prompts that superficially resemble unsafe ones. Fu et al. (2024b) employ a similar practice to improve the robustness of LLMs against malicious queries when processing long text. Another noteworthy work by Lyu et al. (2024) uncovers the crucial role of the prompt templates in preserving safety alignment. The proposed "Pure Tuning, Safe Testing" strategy aims to maintain the model’s safety constraints while enhancing its performance by employing carefully designed prompts. This dual approach helps ensure that the fine-tuning process does not compromise the model’s robustness and safety features.
Furthermore, to address the threat of extraction attacks, Ishibashi & Shimodaira (2024) propose a knowledge sanitization method that fine-tunes LLMs to generate innocuous responses such as "I don’t know" when encountering sensitive data. This approach not only safeguards against the extraction of sensitive information but also maintains the overall performance of LLMs in various tasks. Additionally, Rosati et al. (2024) offer a defense mechanism that effectively prevents LLMs from being maliciously fine-tuned for harmful purposes. This mechanism works by removing information about harmful representations such that it is difficult to recover them during fine-tuning. More importantly, it is able to generalize across different subsets of harm that have not been seen during the defense process, substantially enhancing the robustness of LLMs to harmful fine-tuning attacks.
Pain Point: There is a well-known trade-off in enhancing LLM’s instruction-following capabilities while ensuring they remain safe and reliable. On the other hand, fine-tuning itself can be used both to enhance the safety of LLMs by adding safety-related examples, and to attack LLMs by introducing adversarial examples. The mechanism proposed by Rosati et al. (2024), while seemingly promising, is limited to defending against supervised fine-tuning attacks in LLMs. Also, it requires paired safe and unsafe examples, which makes data collection more expensive and complex.
Future Work: Future work on safety fine-tuning perhaps should strive to achieve a win-win situation for both the safety and utility of LLMs. At the same time, future research should invest in stronger attack settings to emulate worst-case attacks during the fine-tuning process and investigate different types of harm, based on which more effective and comprehensive defense mechanisms need to be developed.
4.4 Post-Safety Alignment
Beyond the regular safety alignment process, post-safety alignment has emerged as a secondary safeguard for LLMs against various potential vulnerabilities. This ongoing process is concerned with seeking solutions from both outside and inside the backbone of LLMs to keep them from generating undesirable behaviors (Zhao et al., 2024c). Accordingly, relevant defense techniques have been developed to mitigate a majority of the safety risks posed by data privacy attacks, prompt injection attacks, and jailbreak attacks.

### v3

@@ 3.4 Energy-Latency Attacks and Potential Threats against LLMs
The energy-latency attacks originate in a broader area of neural networks and their inference processes. The NMTSloth methodology (Feng et al., 2024) introduces a gradient-guided approach to detect efficiency degradation in Neural Machine Translation (NMT) systems. By delaying the appearance of the end-of-sequence (EOS) token through subtle perturbations, NMTSloth demonstrates how altering the output probability distribution can escalate computational demands. Another empirical study (Hong et al., 2020) leverages the adaptive nature of neural networks by introducing subtle perturbations during inference, significantly increasing the model’s inference time.
@@
LLMs, akin to other neural networks, are vulnerable to energy-latency attacks. Methods explored in the aforementioned paper can be similarly adapted for LLMs to heighten computational requirements and prolong response times. According to our research, we have identified that LLMs (LLaMa-series) tend to engage in excessive analysis in some trigger scenarios. For example, when the input data contains numerical values, using the instruction "let’s think step-by-step" prompts the model to prioritize mathematical computations, even if the instruction itself is not directly related to solving math problems. This tendency can lead the model to engage in unnecessary and detailed mathematical analysis, which may not align with the user’s intended query or task.
@@
Although energy-latency attacks have not been extensively explored in the context of LLMs, addressing their potential threats is crucial for ensuring the efficient and reliable operation of these models. While building efficient LLMs has been the focus of extensive research (Yang et al., 2023a; Sanh, 2019; Jiao et al., 2019; Cho, 2014), understanding and mitigating the risks posed by energy-latency attacks remains an important area for future investigation. Enhancing our ability to counteract these attacks will contribute to the development of more resilient and effective LLM systems.
3.5 Limitation and Future Work
These attacks often assume that attackers can freely craft inputs that cause the model to generate excessively long or computationally intensive responses without any restrictions. In real-world deployments, LLMs are typically equipped with input validation and moderation systems that limit the length and complexity of outputs. Service providers can implement maximum input and output lengths, filter out potentially manipulative instructions, and monitor for anomalous usage patterns. Future work could explore how attackers might adapt their strategies in light of robust defenses. This includes developing sophisticated inputs that evade detection or simulate real user behavior to bypass monitoring systems.
Attack Method Vulnerabilities Exploited Attack Surface Attacker Capability Attack Goal Defense Strategy
Attacks on SFT Increased LLM vulnerabilities from SFT and quantization; Overfitting SFT model weights; SFT training data; Fine-tuning APIs White-box or Black-box access; Ability to modify fine-tuning data; Access to fine-tuning APIs Utility loss; Integrity violation Adversarial training; Safety fine-tuning
Attacks on RLHF Increased LLM vulnerabilities from RLHF; Overfitting Model weights; PPO/DPO training data; Reward model training data White-box or Black-box access; Ability to modify PPO/DPO training data or reward model training data Utility loss; Integrity violation Safety fine-tuning; Model merging
Jailbreaks Gap between model capacity and alignment; Intrinsic conflict in LLM objectives Input data; Generation process Black-box attack for prompt-based; White-box for generation-based Integrity violation; Privacy leak Red team defense; Adversarial training; Safety fine-tuning; Content filtering; Inference guidance
Prompt Injection Attacks Model’s over-reliance on input prompts; Prompt parsing weaknesses Input data Black-box attack; Ability to modify input data Integrity violation Red team defense; Content filtering; Adversarial training; Safety fine-tuning
Inference Attacks Model memorization; Overfitting Model outputs Black-box or White-box access; Ability to obtain model outputs Privacy leak Red team defense; Inference guidance; Adversarial training; Safety fine-tuning
Extraction Attacks Model memorization; Overfitting Model outputs Black-box or White-box access; Ability to query the model extensively Privacy leak Adversarial training; Safety fine-tuning
Energy-Latency Attacks Inefficient handling of specific inputs; Lack of resource constraints Model inputs Black-box attack; Ability to craft specific inputs Utility loss Red team defense; Content filtering
Table 1: Overview of attack methods with corresponding vulnerabilities exploited, attack surfaces, attacker capabilities, attack goals, and defense strategies.
@@ 4 Defense
This section explores advanced defense strategies aimed at enhancing the robustness and safety of LLMs. We categorize them into three main groups: exogenous defenses, endogenous defenses, and hybrid defenses. Particularly, exogenous defense strategies prevent models from generating harmful content by setting a series of rules and constraints at the model input or output levels. Specific techniques associated with this paradigm include Red Teaming, Content Filtering, and Inference Guidance. Endogenous defense strategies improve the robustness and safety of LLMs by modifying their internal mechanisms or representations. The typical techniques involve Adversarial Training, Safety Fine-tuning, and Model Merging. Hybrid defense strategies combine the benefits of both endogenous and exogenous defense strategies to provide more comprehensive protection. By comprehensively reviewing these defense strategies, this section aims to provide insights into cutting-edge approaches for defending LLMs, ensuring their secure and effective deployment in real-world applications. Understanding and advancing these defenses are crucial steps toward building trustworthiness and resilience in LLMs, facilitating their responsible integration into various domains of modern society.
@@ 4.1 Red Team Defense
An effective way to discover potential risks in LLMs deployment is to use the Red Team Defense. The Red Team Defense methodology is centered around simulating real-world attack scenarios to uncover potential vulnerabilities and weaknesses in LLMs. The outcomes are used to improve security policies, procedures, and technical defenses. The process includes the following steps:
@@
Challenges: Red teaming can be resource-intensive and requires skilled personnel to effectively mimic sophisticated attack strategies (Xu et al., 2021; Ribeiro et al., 2020; Röttger et al., 2020). The development of red teaming methodologies is still in its early stages, resulting in a lack of comprehensive statistical data on its effectiveness and outcomes.
Recent advancements in leveraging language models for red teaming have introduced automated approaches to simulate test cases and then employ a classifier to detect whether these test cases could lead to harmful outputs from the LLM (Perez et al., 2022). The shift towards automation brings several benefits: firstly, it increases test case diversity and difficulty; secondly, it enables scalable testing across a wide range of scenarios and environments. The work proposed by Ganguli et al. (2022) aimed to build a more efficient AI-assistant interface to collect scale red team data for further analysis. They also studied the scalability of different sizes and types of LLMs under read team attacks, pointing out that rejection sampling and RLHF could build a stronger defense against various attacks.
Besides security and defense purposes, red teaming could help boost LLM performance in different areas as well. In the paper of Buszydlik et al. (2023), the authors use red teaming techniques to simulate different types of mathematical problems and puzzles, then evaluate the performance of LLMs in solving them. While the paper’s focus on mathematical tasks is narrow, its methodology and findings are valuable for understanding the broader impact of red teaming techniques on LLM performance.
@@
4.2 Content Filtering
Content filtering, encompassing input filtering and output filtering, serves as a pivotal exogenous defense strategy designed to safeguard the integrity and appropriateness of interactions with LLMs. By identifying and intercepting harmful input prompts and output contents before dissemination, content filtering ensures that the model’s responses comply with predefined legal, ethical, and contextual guidelines. This process involves understanding and delineating inappropriate content categories, and implementing robust mechanisms to preemptively or reactively filter such content in real time.
Recent advancements in content filtering have witnessed the development of a variety of methods that can be broadly categorized into two mainstreams, i.e., rule-based and learning-based systems. Rule-based filters aim to detect malicious inputs or harmful outputs by applying a series of predefined rules or patterns. For example, observing that queries with adversarial suffixes have exceedingly high perplexity values, Alon & Kamfonas (2023) propose to use perplexity as a metric for detecting adversarial prompts to resist jailbreak attacks. In addition, Jain et al. (2023) introduce a preprocessing approach to modify the expression of the original input by paraphrasing and re-tokenization, thus rendering attacks that are sensitive to the input’s semantic representation ineffective. Based on the finding that adversarial generated prompts are vulnerable to character-level changes, the SmoothLLM method (Robey et al., 2023) randomly perturbs multiple copies of a given input prompt and then aggregates the corresponding predictions to detect adversarial inputs. Moreover, Cao et al. (2023) develop a robust alignment checking function that detects and blocks potentially malicious inputs to defend against alignment-breaking attacks. Kumar et al. (2023) present an “erase-and-check” framework to defend against prompt injection attacks by erasing tokens individually and inspecting the resulting subsequences using a safety filter.
Alternatively, learning-based filters capitalize on the inherent capabilities of powerful models to detect harmful content, leading to adaptable systems that can capture nuanced patterns that rule-based approaches might miss. Specifically, Perspective-API (Google, 2024) and Moderation (OpenAI, 2023) are two popular platforms that prodive a variety of machine learning models to identify potentially harmful or toxic inputs. Beyond that, Chiu et al. (2021) use GPT-3 directly to detect text passages involving sexism and racism. To improve the accuracy of hate speech detection, Goldzycher & Schneider (2022) utilize a BART-large model (Lewis, 2019) to predict different aspects of the input text through multiple hypotheses. Pisano et al. (2023) propose the Bergeron framework, which adopts an auxiliary LLM to protect the primary LLM from incoming attacks, as well as to monitor its output for harmful content. Kim et al. (2023) introduce the Adversarial Prompting Protection (APS), a lightweight model that can detect adversarial prompts effectively.
Challenges: Despite the significant advancements in content filtering, various challenges persist which limit its efficacy and robustness. A primary challenge lies in the adversarial nature of harmful prompts, which continually evolve to evade detection mechanisms. Rule-based filtering, although straightforward, struggles with maintaining relevancy as attackers devise sophisticated inputs that bypass fixed patterns. On the other hand, while learning-based systems show a greater capacity for adaptability, their reliance on large training datasets raises concerns regarding the sufficiency and diversity of such datasets in accurately representing harmful content variations. Moreover, these systems often struggle with balancing sensitivity and specificity, leading to scenarios where benign content is incorrectly flagged or, conversely, harmful content is missed. There is also an ongoing debate regarding the explainability of filtering decisions, with the opaque nature of many machine learning models posing challenges in transparency and acceptance.
Future Work: Looking toward the future, the development of more sophisticated and comprehensive content filtering approaches is imperative. One promising avenue is the integration of hybrid systems that combine rule-based precision with the adaptability of learning-based mechanisms, potentially leveraging real-time feedback to dynamically adjust filtering criteria. Additionally, advancements in transfer learning and few-shot learning methods could enhance the capacity of models to adapt quickly to new types of harmful content with limited input data. As explainability continues to gain importance, research into interpretable machine learning could provide greater clarity and understanding of filtering decisions, fostering trust and ensuring compliance with ethical standards. Furthermore, collaborative efforts that facilitate the sharing of evolving threat patterns and datasets across platforms and organizations might offer a robust framework for collectively bolstering defenses against inappropriate content propagation in large language models.
4.3 Inference Guidance
Inference guidance represents another exogenous defense strategy specifically designed to navigate the inference processes of LLMs to mitigate the generation of harmful content. By actively incorporating additional constraints or guidance during the inference phase, this approach strives to enhance the model’s resilience against jailbreak attacks and ensures that the model’s responses are both safe and useful.
Recent research in inference guidance has yielded a variety of innovative defense methods. One prominent scheme is the use of system prompts, which are essentially an integral part of LLMs due to carrying crucial instructions to guide the models’ behavior (Touvron et al., 2023; Chiang et al., 2023). A well-designed system prompt can elicit the model’s innate security capabilities as much as possible. Specifically, Xie et al. (2023) present a simple yet effective method to defend against various jailbreak attacks by encapsulating the user’s query in a system prompt that reminds ChatGPT to respond responsibly. Furthermore, Wei et al. (2023) explore a strategy called "In-Context Defense" that teaches the LLM to resist jailbreaking by imitating a few examples of refusing harmful queries. Besides, Phute et al. (2023); Zhang et al. (2023c) attempt to highlight safety concerns in system prompts to encourage LLMs to generate responsible outputs.
Beyond prompt-based guidance, another effective paradigm to steer the LLM’s decoding process is directly manipulating the probability of generated tokens. For instance, Zhong et al. (2024) propose a straightforward method based on the principle of contrastive decoding, which aims to boost the probability of desired safe outputs by suppressing undesired outputs. Xu et al. (2024a) develop a safety-aware decoding strategy to protect LLMs from jailbreak attacks. This strategy identifies safety disclaimers and amplifies their token probabilities while attenuating the probabilities of token sequences aligned with jailbreak attacks’ objectives. Additionally, Li et al. (2023b) present RAIN, a search-and-backward framework to control token selection based on the estimated safety of each token. In particular, the search phase is responsible for exploring the potential content that each token may generate and evaluating their safety scores, which are subsequently aggregated in the backward phase to adjust the probabilities for token selection.
Challenges: One primary concern is the trade-off between mitigating harmful content and preserving the naturalness and informativeness of the outputs. Adjusting token probabilities and employing constraining prompts often result in overly conservative responses, which might compromise the quality of information conveyed by the model. Moreover, current methods may introduce inefficiencies during the inference process, as they often require intricate computation or additional data manipulation, thus potentially degrading the model’s performance or response time. Additionally, the susceptibility of LLMs to creative jailbreak methods that evolve alongside the defense mechanisms poses another significant challenge, as attackers continuously develop new ways to bypass existing safety measures. These hurdles underscore the need for more adaptable and context-aware inference guidance techniques that do not excessively compromise model efficiency.
Future Work: Looking forward, the defense strategy of inference guidance for LLMs can benefit from several promising research directions. One potential avenue is the development of dynamic and context-sensitive guidance mechanisms that can adaptively tailor the inference process based on real-time interactions and evolving contexts. Leveraging advances in reinforcement learning and continual learning could facilitate more personalized and contextually aware responses, thereby enhancing the model’s safety without sacrificing versatility. Furthermore, integrating interdisciplinary insights from fields such as ethics, linguistics, and cultural studies could provide a more nuanced understanding of harm in language generation, enabling better alignment of LLMs with diverse societal norms and values. Finally, the emergence of collaborative frameworks that combine user feedback with automated moderation tools could offer a more robust and resilient defense against ever-evolving threats, effectively balancing safety considerations with the desired efficacy and adaptability of LLMs.
@@ 4.4 Adversarial Training
Challenges: Given the vast input space, comprehensively identifying failure modes in LLMs is challenging and resource-intensive. Besides, defenders often focus on robustness-related failure modes, such as those involving Lp-norm attacks. However, attackers may employ more subtle methods, including Trojans and jailbreaks, which are harder to detect. And adversarial training often causes a trade-off between robustness and performance on clean data (Min et al., 2021; Raghunathan et al., 2020). Furthermore, LLMs’ responses to adversarial training are less effective than expected. They might memorize adversarial samples used in training rather than developing generalized defenses. And they struggle to modify pre-train gained knowledge (Prakash et al., 2024)
@@
Recent studies have introduced new approaches to address these challenges. Liu et al. (2020) find that adversarial pre-training using ALUM (Adversarial Learning with Unlabeled Model) leads to improvements in both generalization and robustness across a range of natural language processing tasks. The use of virtual adversarial training objectives (Miyato et al., 2018) allows ALUM to smooth the embedding space of the model and balance standard error with robust error effectively. This capability suggests a promising direction for adversarial training in LLMs, integrating robust defense strategies without compromising model performance. To further address the computational inefficiencies of traditional adversarial training methods, Xhonneux et al. (2024) introduce two novel algorithms that perform adversarial attacks within the continuous embedding space of LLMs. These algorithms—CAT (Continuous Adversarial Training) and CAPO (Continuous Adversarial Perturbation Optimization)—significantly enhance model robustness against discrete adversarial attacks while maintaining utility. By operating in the continuous embedding space, these algorithms reduce computational overhead and provide scalable solutions for adversarial training. To improve the generalization of adversarial training, Casper et al. (2024) introduce latent adversarial training (LAT) to learn more robust representations by incorporating latent adversarial examples during training. Huang et al. (2024e) propose a technique (Vaccine) that aims to produce invariant hidden embeddings by progressively adding crafted perturbations in the alignment phase. This enables the embeddings to withstand harmful perturbations from un-sanitized user data during fine-tuning. Both LAT and Vaccine operate at the hidden embedding level, distinguishing them from traditional methods that operate at the input level, which helps to address potential shifts in data distribution that may occur between the development and deployment phases in real-world applications. By introducing carefully designed perturbations in the latent space, they enhance the model’s robustness against such shifts, which could introduce Trojans, jailbreaks, and other unforeseen challenges.
@@
Future Work: Building on the insights from Casper et al. (2024) and Huang et al. (2024e), future research could explore the implementation of latent adversarial training as an alternative to traditional adversarial training during the fine-tuning stage. Additionally, incorporating adversarial training into the pre-training phase of model development could further enhance robustness against attacks. These approaches aim to improve the model’s resilience by addressing vulnerabilities early in the training process and better equipping the model to handle adversarial inputs.
@@ 4.5 Safety Fine-Tuning
Fine-tuning has gained great popularity among end-users as a core technique for customizing pre-trained LLMs to various downstream tasks. However, recent research has revealed potential safety risks associated with exercising this privilege. Qi et al. (2023) find that the safety alignment of LLMs can be compromised by fine-tuning with only a few adversarially designed training examples. In addition, they further suggest that even without malicious intent, simply fine-tuning with benign and commonly used datasets may also inadvertently degrade the safety alignment of LLMs. To investigate the reason behind this, Qi et al. (2024) identify the existence of the “shallow alignment” issue, i.e., safety alignment often takes shortcuts by adapting a model’s generative distribution primarily over only its very first few output tokens. As a result, they propose to use data augmentation and constrained optimization objectives to make the alignment deeper than just a few tokens.
To address the vulnerabilities introduced by fine-tuning, a natural solution one would expect is to incorporate safety-related examples during the fine-tuning stage. Bianchi et al. (2023) and Zhao et al. (2023) verify the effectiveness of this approach. They suggest that including a small number of safety examples in the fine-tuning process can significantly enhance the safety of LLMs while not degrading their usefulness. Further, Zong et al. (2024) obtain the same conclusion by including safety-related examples during the fine-tuning process of vision-LLM. Accordingly, Fu et al. (2024) employ a similar practice to improve the robustness of LLMs against malicious queries when processing long text. Unfortunately, Huang et al. (2024a) find that both alignment stage and fine-tuning stage defenses fail when specific training hyper-parameters are chosen, such as a large learning rate or a large number of training epochs. Thus, they introduce a one-shot pruning stage after harmful fine-tuning to remove the harmful weights responsible for generating harmful content. In addition, by exploring different directions of perturbing model weights, Peng et al. (2024) discover a "safety basin" in the model parameter space, where randomly perturbing model weights maintain the safety level of the original aligned model in its local neighborhood.
To mitigate the jailbreaks effect when fine-tuning on datasets mixed with harmful data, Huang et al. (2024d) propose to optimize the alignment and user datasets independently by separating the states in the fine-tuning stage. Meanwhile, they introduce a proximal term to constraint the drift of each state for solving the problem of unstable convergence and degraded alignment performance during the optimization process. Toward the same goal, Liu et al. (2024c) present a data curation framework to mitigate jailbreaks while preserving model usefulness. This framework operates under the assumption of no prior knowledge of the attack specifics and is designed to mitigate model harmful responses caused by toxic data during either the pre-training or fine-tuning stages. To further reduce safety fine-tuning sample complexity against jailbreaks effects, Wang et al. (2024) propose a method inspired by the efficiency of backdoor attacks, which require only a small amount of poisoned data to achieve the attack target. The method, called Backdoor Enhanced Safety Alignment, constructs prefixed safety examples by integrating a secret prompt, acting as a "backdoor trigger," that is prefixed to safety examples. This is particularly important for practical applications where the availability of safety examples may be limited.
To mitigate the risk of fine-tuning while maintaining accuracy, Huang et al. (2024b) propose to solve a perturbation minimization problem by constraining the gap of harmful loss before and after a simulated perturbation. The proposed problems are solved by an iterative gradient method named Booster. Moreover, Hsu et al. (2024) propose Safe LoRA, which aims to address the safety risks associated with fine-tuning LLM using the LoRA approach. The authors propose a simple one-liner patch to the original LoRA implementation by introducing the projection of LoRA weights from selected layers to the safety-aligned subspace, effectively reducing the safety risks in LLM fine-tuning while maintaining utility. Besides, Yi et al. (2024b) introduce a safety realignment framework for LLM based on subspace-oriented model fusion (SOMF), aiming to combine the safeguard capabilities of initially aligned models and current fine-tuned models into a realigned model. The approach begins by disentangling all task vectors from the weights of each fine-tuned model, identifying safety-related regions within these vectors by subspace masking techniques, and finally exploring the fusion of the initial safely aligned LLM with all task vectors based on the identified safety subspace. The framework is validated to satisfy the safety requirements of a single fine-tuned model as well as multiple models during their fusion. Hence, it can be naturally extended to the scenario of simultaneously restoring safety during the fusion of multiple fine-tuned models.
To comprehensively understand the safety risks associated with fine-tuning, Leong et al. (2024) analyze the mechanisms of different fine-tuning attacks. They break down the safeguarding process of an LLM when encountering harmful instructions into three stages: recognizing harmful instructions, generating an initial refusing tone, and completing the refusal response. They then analyze how two representative types of attack approaches, Explicit Harmful Attack (EHA) and Identity-Shifting Attack (ISA), influence each stage of this safeguarding process. And they suggest that diverse defense mechanisms are needed to effectively cope with these attacks. A noteworthy work of Lyu et al. (2024) uncovers the crucial role of the prompt templates in preserving safety alignment. The proposed "Pure Tuning, Safe Testing" strategy aims to maintain the model’s safety constraints while enhancing its performance by employing carefully designed prompts. This dual approach helps ensure that the fine-tuning process does not compromise the model’s robustness and safety features. Furthermore, to address the threat of extraction attacks, Ishibashi & Shimodaira (2023) propose a knowledge sanitization method that fine-tunes LLMs to generate innocuous responses such as "I don’t know" when encountering sensitive data. This approach not only safeguards against the extraction of sensitive information but also maintains the overall performance of LLMs in various tasks.